# Peeling metric spaces of strict negative type

### v1

@@ 1 Introduction
Many problems in data science and machine learning can be distilled to identifying outliers boukerche2020outlier , anomalies chandola2009anomaly , and diverse subsets leinster2016maximizing ; mahabadi2023core . A vast and almost totally disconnected literature that we shall not attempt to capture with additional references is devoted to these problems.
This paper details a unified natural interpretation of, and framework for, solving these problems in a broad class of situations. Specifically, so-called strict negative type finite metric spaces (including, but not limited to, finite subsets of Euclidean space) admit a natural notion of outliers or boundary elements that we call a peel and that simultaneously maximizes a natural measure of diversity leinster2021entropy . The notion of a peel involves no ambiguity (e.g., free parameters) and a peel can be computed by solving a finite (and in practice, short) sequence of linear equations, as detailed in Algorithm 1 below. After proving the correctness of Algorithm 1, we detail its applicability to mitigating hallucinations in large language models. We then discuss product metrics, with an eye towards computing outlying/diverse sequences or paths, including a detailed real-world example. Appendices contain proofs, discussions of extensions, and auxiliary experimental results.
@@ 2 Weightings, magnitude, and diversity
We say that $d$ is negative type if $x^{T}dx\leq 0$ for $1^{T}x=0$ and $x^{T}x=1$ (equivalently to this last, $x\neq 0$). If the inequality is strict, we say that $d$ is strict negative type: this entails that $Z$ is positive semidefinite for all $t>0$. Important examples of negative type metrics on finite spaces are finite subsets of Euclidean space with the $L^{1}$ or $L^{2}$ distances, finite subsets of spheres with the geodesic distance, finite subsets of hyperbolic space, and ultrametrics (i.e., metrics satisfying $d(x,z)\leq\max\{d(x,y),d(y,z)\}$) on finite spaces. However, not all of these are strict negative type: e.g., spheres with antipodal points are not strict negative type hjorth1998finite .
A weighting $w$ is a solution to $Zw=1$, where $1$ indicates a vector of all ones. If $Z$ has a weighting $w$, then its magnitude is $\text{Mag}(Z):=\sum_{j}w_{j}$. If $d$ is negative type, then $Z$ is positive definite, so it has a unique weighting. It turns out that weightings are excellent scale-dependent boundary or outlier detectors in Euclidean space willerton2009heuristic ; bunch2020practical ; huntsman2023diversity : in fact, behavior evocative of boundary detection applies more generally huntsman2023magnitude . A technical explanation of the Euclidean boundary-detecting behavior draws on the notion of Bessel capacities meckes2015magnitude .
@@
Fairly recent mathematical developments have clarified the role that magnitude and weightings play in maximizing a general and axiomatically supported notion of diversity leinster2016maximizing ; leinster2021entropy . Specifically, the diversity of order $q$ for a probability distribution $p$ and similarity matrix $Z$ is
@@
If $Z$ is symmetric, positive definite, and has a unique positive weighting $w$, then for all $q$, $w$ is proportional to the diversity-maximizing distribution leinster2016maximizing .
The situation described by Theorem 1 reduces diversity maximization to a standard linear algebra problem while simultaneously removing any ambiguity regarding the parameter $q$. It is possible to efficiently compute a “cutoff scale” huntsman2023diversity such that we can optimally enforce this desirable situation for similarity matrices of the form $Z=\exp[-td]$. However, in practice this scale is often quite large, and the resulting weighting will have many components with values close to unity, degrading the utility of this construction. It is frequently desirable to work in the limit $t\downarrow 0$: for example, in Figure 1, this limit successfully identifies one point as an outlier. We turn to this limit in the sequel.
@@ 3 The peeling theorem
The quantity $p^{T}dp$ is called the quadratic entropy of $d$: it is convex if $d$ is strict negative type. (For details, see Theorem 4.3 of rao1984convexity and Proposition 5.20 of devriendt2022graph as well as leinster2012measuring ; meckes2013positive ; leinster2016maximizing ; leinster2021entropy .) Therefore if $d$ is strict negative type, (4) can be efficiently maximized over any sufficiently simple polytope via quadratic programming. However, in §3.1 we will give a more practical (i.e., much faster and more sparsity-accurate) algorithm for maximizing the quadratic entropy of strict negative type metrics.
@@ 3.1 Maximizing quadratic entropy of strict negative type metrics
Translated into our context, Proposition 5.20 of devriendt2022graph states that if $d$ is strict negative type, then
@@
The following theorem (with proof in §A.1) generalizes Theorem 5.23 of devriendt2022graph and addresses the $t\downarrow 0$ limit of an algorithm successively sketched and fully described in preprint versions of huntsman2023diversity and huntsman2023quality but omitted from the published versions.
@@ 3.2 Iterated peeling of text embeddings
We selected 150 named RGB color codes from the large-scale color survey xkcd ; lo2024exploring by restricting consideration to colors with a single word in their name, and then further restricting by human judgment to get a desired number while trying to avoid ambiguity. We then fed prompts of the form
@@
to gpt-4o, where the placeholder is for a color name. We embedded prompts and responses using voyage-3.5  [^1] See https://blog.voyageai.com/2025/05/20/voyage-3-5/. ModernBERT warner2024smarter produced visually inferior embeddings (not shown, but see lo2024exploring ). and repeatedly peeled the results using spherical distance of normalizations, as shown in Figures 3-6. Appendix §C shows an example along the same lines with all 150 colors at once.
@@
As another example informed by a survey of numerical score assignments for sentiment words in yougov , we fed prompts of the form
@@ 4.1 $L^{p}$ products of strict negative type metrics
where $J$ is a matrix of all ones deza2009encyclopedia . That is,
@@
While positive definite spaces are closed under $L^{1}$ products, the same is not true for $L^{p}$ products for any $p>1$ meckes2013positive . This suggests that any attempt to prove that $L^{p}$ products of strict negative type spaces are (or are not) also strict negative type cannot be totally trivial.
@@ 4.2 An application to path diversity
Most existing quality-diversity algorithms are not naturally suited for path spaces, even when they only require the existence of a suitable dissimilarity huntsman2023quality . One reason is that the “correct” notion of dissimilarity between variable-length paths is usually a form of edit distance with insertions and deletions. Such distances are notoriously tricky to handle, particularly with respect to considerations of magnitude and diversity: for example, an embedding of edit distance on $\{0,1\}^{n}$ into $L^{1}$ requires distortion $\Omega(\log n)$ krauthgamer2009improved . Another reason is that path spaces scale exponentially, and computing diversity or a proxy thereof over a path space is intractable without sacrifices in some direction.
Consider a space of fixed-length paths of the form $(v_{1},\dots,v_{L})\in\prod_{\ell=1}^{L}V_{\ell}$ for $L>1$, and suppose that we have a text description associated to each $V_{\ell}$. It is generally straightforward to produce associated embeddings $X_{\ell}$, though it is also generally infeasible to produce embeddings for the entire path space $\prod_{\ell}V_{\ell}$. The usual metric for each $X_{\ell}$ is geodesic (i.e., cosine) distance on the sphere (via normalization), which is negative type and also strict negative type unless $X_{\ell}$ contains antipodes hjorth1998finite . It is not particularly abusive to claim that each $X_{\ell}$ is almost surely strict negative type. By Theorem 3, $\prod_{\ell}X_{\ell}$ is almost surely strict negative type under the $L^{2}$ product distance.
@@ 4.2.1 Example
Let $V_{\ell}=V$ given by the 80 largest US cities as listed in wiki:List_of_United_States_cities_by_population in July 2025. We construct text features for each city using their coordinates and Köppen-Geiger classifications produced by the Python package kgcpy rubel2017climate . The text features are templated like:
@@
Next, we form a directed acyclic graph (DAG) on $V$ with arcs $(v,v^{\prime})$ only for city pairs such that the Euclidean vector from (the planar longitude/latitude coordinates of) $v$ to $v^{\prime}$ has a positive inner product with the Euclidean vector from New York (NY) to Los Angeles (LA). We then restrict this DAG to the vertices/cities with at least one incident arc. By construction, this DAG has a single source at NY and a single target at LA. We then consider the 500 geographically shortest paths from NY to LA in this DAG that have two intermediate stops.  [^2] This example was inspired by the Cannonball Run Challenge needham1981cannonball ; yates2003cannonball .
@@
In particular, the path from NY to Lincoln, Nebraska to Aurora, Colorado to LA explicitly involves traversing Dfa (hot-summer humid continental) and BSk (cold semi-arid) Köppen-Geiger climates in the Great Plains and approaching the Rocky Mountains, respectively.  [^3] The large discrepancy in geography and climate between Lincoln, Nebraska and the Rocky Mountains is a significant plot point in the film farrelly1994dumb .
@@ A.1 Proof of Theorem 2
First, we show that condition ii) is maintained throughout the while loop of Algorithm 1. Note that the condition trivially holds at initialization; also note that any nonempty principal submatrix $d_{\mathcal{J},\mathcal{J}}$ of $d$ is strict negative type and hence invertible. Block partitioning $d$ and $p$ respectively as
$d=\begin{pmatrix}d_{\mathcal{J},\mathcal{J}}&d_{\mathcal{J},\mathcal{J}^{c}}\\ d_{\mathcal{J}^{c},\mathcal{J}}&d_{\mathcal{J}^{c},\mathcal{J}^{c}}\end{pmatrix}$;
$p=p_{\mathcal{J}}\oplus 0_{\mathcal{J}^{c}}=\frac{d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}{1_{\mathcal{J}}^{T}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}\oplus 0_{\mathcal{J}^{c}},$
where $\mathcal{J}^{c}$ denotes the set complement $n]\backslash\mathcal{J}$, we obtain
$dp=\frac{1_{\mathcal{J}}\oplus d_{\mathcal{J}^{c},\mathcal{J}}p_{\mathcal{J}}}{1_{\mathcal{J}}^{T}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}.$
Thus for $j\in\mathcal{J}$
$e_{j}^{T}dp=\frac{1}{1_{\mathcal{J}}^{T}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}},$
and for $k\in\mathcal{J}^{c}$
$e_{k}^{T}dp=\frac{d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}{1_{\mathcal{J}}^{T}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}.$
Now the case $k\in\mathcal{J}$ for condition ii) is trivial, so condition ii) is equivalent to
$1\geq d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}$
for all $k\in\mathcal{J}^{c}$. We rewrite the right hand side until it is obvious that an equivalent of condition ii) is satisfied:
$d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}$ $=\sum_{j^{\prime}\in\mathcal{J}}\sum_{j\in\mathcal{J}}d_{kj}(d^{-1})_{jj^{\prime}}$
$=\sum_{j^{\prime}\in\mathcal{J}}(dd^{-1})_{kj^{\prime}}$
$=\sum_{j^{\prime}\in\mathcal{J}}\delta_{kj^{\prime}}$
$=0$
where the last equality is because $k\in\mathcal{J}^{c}$.
Algorithm 1 halts once condition i) holds. The computational complexity bound holds because the while loop takes at most $n$ iterations. ∎

### v2

@@ 1 Introduction
Many problems in data science and machine learning can be distilled to identifying outliers Boukerche et al. (2020), anomalies Chandola et al. (2009), and diverse subsets Leinster and Meckes (2016); Mahabadi and Trajanovski (2023). A vast and almost totally disconnected literature that we shall not attempt to capture with additional references is devoted to these problems.
This paper details a unified natural interpretation of, and framework for, solving these problems in a broad class of situations. Specifically, so-called strict negative type finite metric spaces (including, but not limited to, finite subsets of Euclidean space) admit a natural notion of outliers or boundary elements that we call a peel and that simultaneously maximizes a natural measure of diversity Leinster (2021). The notion of a peel involves no ambiguity (e.g., free parameters) and a peel can be computed by solving a finite (and in practice, short) sequence of linear equations, as detailed in Algorithm 1 below. After proving the correctness of Algorithm 1, we detail its applicability to mitigating hallucinations in large language models. We then discuss product metrics, with an eye towards computing outlying/diverse sequences or paths, including a detailed real-world example. Appendices contain proofs, discussions of extensions, and auxiliary experimental results.
@@ 2 Weightings, magnitude, and diversity
We say that $d$ is negative type if $x^{T}dx\leq 0$ for $1^{T}x=0$ and $x^{T}x=1$ (equivalently to this last, $x\neq 0$). If the inequality is strict, we say that $d$ is strict negative type: this entails that $Z$ is positive semidefinite for all $t>0$. Important examples of negative type metrics on finite spaces are finite subsets of Euclidean space with the $L^{1}$ or $L^{2}$ distances, finite subsets of spheres with the geodesic distance, finite subsets of hyperbolic space, and ultrametrics (i.e., metrics satisfying $d(x,z)\leq\max\{d(x,y),d(y,z)\}$) on finite spaces. However, not all of these are strict negative type: e.g., spheres with antipodal points are not strict negative type Hjorth et al. (1998).
A weighting $w$ is a solution to $Zw=1$, where $1$ indicates a vector of all ones. If $Z$ has a weighting $w$, then its magnitude is $\text{Mag}(Z):=\sum_{j}w_{j}$. If $d$ is negative type, then $Z$ is positive definite, so it has a unique weighting. It turns out that weightings are excellent scale-dependent boundary or outlier detectors in Euclidean space Willerton (2009); Bunch et al. (2020); Huntsman (2023a): in fact, behavior evocative of boundary detection applies more generally Huntsman (2023b). A technical explanation of the Euclidean boundary-detecting behavior draws on the notion of Bessel capacities Meckes (2015).
@@
Fairly recent mathematical developments have clarified the role that magnitude and weightings play in maximizing a general and axiomatically supported notion of diversity Leinster and Meckes (2016); Leinster (2021). Specifically, the diversity of order $q$ for a probability distribution $p$ and similarity matrix $Z$ is
@@
If $Z$ is symmetric, positive definite, and has a unique positive weighting $w$, then for all $q$, $w$ is proportional to the diversity-maximizing distribution Leinster and Meckes (2016).
The situation described by Theorem 1 reduces diversity maximization to a standard linear algebra problem while simultaneously removing any ambiguity regarding the parameter $q$. It is possible to efficiently compute a “cutoff scale” Huntsman (2023a) such that we can optimally enforce this desirable situation for similarity matrices of the form $Z=\exp[-td]$. However, in practice this scale is often quite large, and the resulting weighting will have many components with values close to unity, degrading the utility of this construction. It is frequently desirable to work in the limit $t\downarrow 0$: for example, in Figure 1, this limit successfully identifies one point as an outlier. We turn to this limit in the sequel.
@@ 3 The peeling theorem
The quantity $p^{T}dp$ is called the quadratic entropy of $d$: it is convex if $d$ is strict negative type. (For details, see Theorem 4.3 of Rao (1984) and Proposition 5.20 of Devriendt (2022) as well as Leinster and Cobbold (2012); Meckes (2013); Leinster and Meckes (2016); Leinster (2021).) Therefore if $d$ is strict negative type, (4) can be efficiently maximized over any sufficiently simple polytope via quadratic programming. However, in §3.1 we will give a more practical (i.e., much faster and more sparsity-accurate) algorithm for maximizing the quadratic entropy of strict negative type metrics.
@@ 3.1 Maximizing quadratic entropy of strict negative type metrics
Translated into our context, Proposition 5.20 of Devriendt (2022) states that if $d$ is strict negative type, then
@@
The following theorem (with proof in §A.1) generalizes Theorem 5.23 of Devriendt (2022) and addresses the $t\downarrow 0$ limit of an algorithm successively sketched and fully described in preprint versions of Huntsman (2023a) and Huntsman (2023c) but omitted from the published versions.
@@ 3.2 Iterated peeling of text embeddings
We selected 150 named RGB color codes from the large-scale color survey Munroe (2010); Lo and Hsiao (2024) by restricting consideration to colors with a single word in their name, and then further restricting by human judgment to get a desired number while trying to avoid ambiguity. We then fed prompts of the form
@@
to gpt-4o, where the placeholder is for a color name. We embedded prompts and responses using voyage-3.5  [^1] See https://blog.voyageai.com/2025/05/20/voyage-3-5/. ModernBERT Warner et al. (2024) produced visually inferior embeddings (not shown, but see Lo and Hsiao (2024)). and repeatedly peeled the results using spherical distance of normalizations, as shown in Figures 3-6. Appendix §C shows an example along the same lines with all 150 colors at once.
@@
As another example informed by a survey of numerical score assignments for sentiment words in YouGov (2018), we fed prompts of the form
@@ 4.1 $L^{p}$ products of strict negative type metrics
where $J$ is a matrix of all ones Deza and Deza (2009). That is,
@@
While positive definite spaces are closed under $L^{1}$ products, the same is not true for $L^{p}$ products for any $p>1$ Meckes (2013). This suggests that any attempt to prove that $L^{p}$ products of strict negative type spaces are (or are not) also strict negative type cannot be totally trivial.
@@ 4.2 An application to path diversity
Most existing quality-diversity algorithms are not naturally suited for path spaces, even when they only require the existence of a suitable dissimilarity Huntsman (2023c). One reason is that the “correct” notion of dissimilarity between variable-length paths is usually a form of edit distance with insertions and deletions. Such distances are notoriously tricky to handle, particularly with respect to considerations of magnitude and diversity: for example, an embedding of edit distance on $\{0,1\}^{n}$ into $L^{1}$ requires distortion $\Omega(\log n)$ Krauthgamer and Rabani (2009). Another reason is that path spaces scale exponentially, and computing diversity or a proxy thereof over a path space is intractable without sacrifices in some direction.
Consider a space of fixed-length paths of the form $(v_{1},\dots,v_{L})\in\prod_{\ell=1}^{L}V_{\ell}$ for $L>1$, and suppose that we have a text description associated to each $V_{\ell}$. It is generally straightforward to produce associated embeddings $X_{\ell}$, though it is also generally infeasible to produce embeddings for the entire path space $\prod_{\ell}V_{\ell}$. The usual metric for each $X_{\ell}$ is geodesic (i.e., cosine) distance on the sphere (via normalization), which is negative type and also strict negative type unless $X_{\ell}$ contains antipodes Hjorth et al. (1998). It is not particularly abusive to claim that each $X_{\ell}$ is almost surely strict negative type. By Theorem 3, $\prod_{\ell}X_{\ell}$ is almost surely strict negative type under the $L^{2}$ product distance.
@@ 4.2.1 Example
Let $V_{\ell}=V$ given by the 80 largest US cities as listed in Wikipedia (2025) in July 2025. We construct text features for each city using their coordinates and Köppen-Geiger classifications produced by the Python package kgcpy Rubel et al. (2017). The text features are templated like:
@@
Next, we form a directed acyclic graph (DAG) on $V$ with arcs $(v,v^{\prime})$ only for city pairs such that the Euclidean vector from (the planar longitude/latitude coordinates of) $v$ to $v^{\prime}$ has a positive inner product with the Euclidean vector from New York (NY) to Los Angeles (LA). We then restrict this DAG to the vertices/cities with at least one incident arc. By construction, this DAG has a single source at NY and a single target at LA. We then consider the 500 geographically shortest paths from NY to LA in this DAG that have two intermediate stops.  [^2] This example was inspired by the Cannonball Run Challenge Needham (1981); Yates (2003).
@@
In particular, the path from NY to Lincoln, Nebraska to Aurora, Colorado to LA explicitly involves traversing Dfa (hot-summer humid continental) and BSk (cold semi-arid) Köppen-Geiger climates in the Great Plains and approaching the Rocky Mountains, respectively.  [^3] The large discrepancy in geography and climate between Lincoln, Nebraska and the Rocky Mountains is a significant plot point in the film Farrelly (1994).
@@ A.1 Proof of Theorem 2
 [^4] This is a substantial revision of an earlier proof that had an error in the second claimed equality in a chain of equalities near the end. This revision draws on a session with Claude Opus at https://claude.ai/share/65e7edc0-c0c3-4c0b-915c-36179b719dfe.
Algorithm 1 terminates as soon as condition i) holds. Each iteration of the while loop requires $O(n^{\omega})$ operations and there are at most $n$ iterations, which conditionally establishes the computational complexity bound. Since the condition $e_{k}^{T}dp\geq e_{j}^{T}dp$ for $j\in\mathcal{J}$ trivially holds at initialization, it therefore suffices to show that this condition is maintained throughout the while loop as $\mathcal{J}$ and $p$ are updated.
For $j\in\mathcal{J}$, we have that (abusively writing $e_{j}$ for the $j$th standard basis vector in both $\mathbb{R}^{n}$ and $\mathbb{R}^{|\mathcal{J}|}$)
$e_{j}^{T}dp=\frac{e_{j}^{T}d_{\mathcal{J},\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}{1_{\mathcal{J}}^{T}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}=\frac{1}{1_{\mathcal{J}}^{T}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}=:c.$
(Note that any nonempty principal submatrix $d_{\mathcal{J},\mathcal{J}}$ of $d$ is strict negative type and hence invertible.) Meanwhile for $k\in\mathcal{J}^{c}\equiv[n]\backslash\mathcal{J}$,
$e_{k}^{T}dp=\frac{d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}{1_{\mathcal{J}}^{T}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}}=c\cdot\alpha_{k},$
where $\alpha_{k}:=d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}$. We need to show that $\alpha_{k}\leq 1$ for $k\in\mathcal{J}^{c}$.
Accordingly, write $\mathcal{J}^{\prime}:=\mathcal{J}\cup\{k\}$ and
$d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}=\begin{pmatrix}d_{\mathcal{J},\mathcal{J}}&d_{\mathcal{J},k}\\ d_{k,\mathcal{J}}&0\end{pmatrix}$
so that the Schur complement is $S:=d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}/d_{\mathcal{J},\mathcal{J}}=-d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}d_{\mathcal{J},k}$. Meanwhile, the Schur complement formula yields
$S=\frac{\det d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}}{\det d_{\mathcal{J},\mathcal{J}}}.$
Since both $d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}$ and $d_{\mathcal{J},\mathcal{J}}$ are strict negative type, each has exactly one positive eigenvalue, with the remainder negative. Therefore $S<0$.
Block inverting, we have that
$d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}^{-1}=\begin{pmatrix}d_{\mathcal{J},\mathcal{J}}^{-1}+d_{\mathcal{J},\mathcal{J}}^{-1}d_{\mathcal{J},k}S^{-1}d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}&-d_{\mathcal{J},\mathcal{J}}^{-1}d_{\mathcal{J},k}S^{-1}\\ -S^{-1}d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}&S^{-1}\end{pmatrix}$.
This yields that
$e_{k}^{T}d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}^{-1}1_{\mathcal{J}^{\prime}}=-S^{-1}d_{k,\mathcal{J}}d_{\mathcal{J},\mathcal{J}}^{-1}1_{\mathcal{J}}+S^{-1}=S^{-1}(1-\alpha_{k}).$
Since $S^{-1}<0$, we now have $e_{k}^{T}d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}^{-1}1_{\mathcal{J}^{\prime}}<0\iff\alpha_{k}<1$. Suppose instead that $\alpha_{k}\geq 1$: then $e_{k}^{T}d_{\mathcal{J}^{\prime},\mathcal{J}^{\prime}}^{-1}1_{\mathcal{J}^{\prime}}\geq 0$.
Applying this observation repeatedly for each element removed from a preceding instance of $\mathcal{J}$ in the while loop shows that $k$ would not have been removed from $\mathcal{J}$ in the preceding iteration, contradicting the assumption $k\in\mathcal{J}^{c}$. ∎